OpenStudy (anonymous):

For 2x - 3y = -7, when x = 4, What is the value of y? Choose one answer. a. b. c. -5 d. 5

OpenStudy (anonymous):

So to solve this, lets let x = 4 by replacing x with 4 everywhere in your equation: \[2x-3y=-7 \Rightarrow 2(4)-3y=-7 \Rightarrow 8-3y=-7 \Rightarrow -3y = -15 \Rightarrow y = 5\]

OpenStudy (anonymous):

"lets let".....sounds retarded >.<

OpenStudy (anonymous):

if there are any steps in there that seem confusing, or like witchcraft, let me know, i'll explain :)